OpenStudy (anonymous):

evaluate the series 1/3+1/9+1/27+1/81 usin the formula a1(1-rn)/1-r

OpenStudy (jdoe0001):

so what do you think is the common ratio?

OpenStudy (texaschic101):

this is a geometric series.....to find the common ratio you need to divide the second number by the first number

OpenStudy (ybarrap):

\[\sum_{i=0}^{N}a ^{i}=\frac{ 1-a ^{N+1} }{ 1-a }\] But you don't start at zero, so you'll need to subtract 1 from your answer. Here a = 1/3, N = ?